A planet has a radius of 3500 km and a surface gravity of 3.8 m/s2. what is the mass of the planet

1
Expert's answer
2022-01-30T15:14:57-0500

"g=\frac{GM}{R^2},\implies M=\frac{gR^2}G,"

"M=6.98\cdot 10^{23}~kg."